OpenStudy (anonymous):

3/4(x + 3)= to what ??????

OpenStudy (amistre64):

what are you trying to solve for?

OpenStudy (anonymous):

the answere to the ?

OpenStudy (amistre64):

what you have written makes no sense mathmatically. If we knew what the answer was spose to look like, do you have any options, that would be helpful. If you have any instructions attached to the question, that would help to narrow down what a suitable solution would be as well.

OpenStudy (anonymous):

find x

OpenStudy (amistre64):

x isnt lost. It is a variable that can be used to determine the output of the equation for any given value of x. There is no 'one' solution to x.

OpenStudy (amistre64):

when x = 0; 3/4((0) + 3) = 9/4 when x = 1; 3/4((1) + 3) = 3 when x = any number; 3/4(n + 3) = 3(n+3)/4

OpenStudy (anonymous):

????? well how about idk what you just said its just a problem to solve

OpenStudy (anonymous):

gettin all technical =)

OpenStudy (amistre64):

:) its only solvable when there are rules applied to define it in a sensible fashion; as is it just creates a line on paper to defines all of its solutions

OpenStudy (anonymous):

so do you usually talk all technical

OpenStudy (amistre64):

usually, but if anything, this is the best solution as is ;)

OpenStudy (anonymous):

i agree with you=) were you from

OpenStudy (amistre64):

im from long ago and not to far away. But lets not turn this into something too personal for a public forum please

OpenStudy (anonymous):

oh ok -_-

OpenStudy (amistre64):

if your simply bored, see what going on in the group chat; or maybe read a book :)

OpenStudy (anonymous):

i am well can i ask you how old you are

OpenStudy (amistre64):

at my age, i simply lost track.... i was borned about 1972

OpenStudy (anonymous):

you cant be that old youll be dead